OpenStudy (anonymous):

The price (in dollars) p and the quantity demanded q are related by the equation: p^2+2q^2=1100.If R is revenue, dR/dt can be expressed by the following equation: dR/dt=A*dp/dt, where A is a function of just q. A = ? Find dR/dt when q=15 and dp/dt=5. dR/dt = The price (in dollars) p and the quantity demanded q are related by the equation: p^2+2q^2=1100.If R is revenue, dR/dt can be expressed by the following equation: dR/dt=A*dp/dt, where A is a function of just q. A = ? Find dR/dt when q=15 and dp/dt=5. dR/dt = @Mathematics

OpenStudy (anonymous):

I thought \(R=pq\) so then \[ \frac{dR}{dt} = \frac{dp}{dt}q+p\frac{dq}{dt} \]